# Compositional Deep Probabilistic Models of DNA Encoded Libraries

### v1

@@ 4.6 DEL-Compose performs competitively even in low-resource regimes
Figure 7: Performance of models using factorized representations vs using full molecule representations. Each model is trained with a different % of the data heldout.
@@ 4.7 DEL-Compose facilitates structured interpretation of the data
In Figure 9, we plot the attention distribution over the set of synthon representations of DEL-Compose. The top set of plots aggregate known pharmacophores, while the bottom set of plots delineate individual substructure sets. For CA-IX, we see that the attention probabilities are primarily focused on di-synthon $x_{bc}$, while for HRP, the attention probabilities are primarily distributed on the mono-synthon $x_{b}$. The sulfonamides are on the C position for CA-IX, while the electrophiles are on the B position for HRP, indicating that both models chose the highest weight to be placed on synthon embeddings incorporating the correct synthon position. Moreover, the weights on the important mono/di-synthons are higher for the molecules with the known pharmacophores versus other molecules in the library. Interestingly, the model for CA-IX chooses the di-synthon $x_{bc}$, instead of just the mono-synthon at the C position, which might indicate enrichment of certain di-synthons.
@@ 6 Data and Software Availability
Software: we use RDKit (version 2020.09.125 to parse our molecules and generate Morgan Fingerprints. The details to replicate our model is in the Methods section, and further training details are provided in the Experiments section.

### v2

@@ 4.6 DEL-Compose performs competitively even in low-resource regimes
Figure 6: Performance of models using factorized representations vs using full molecule representations. Each model is trained on the same random splits with a different % of the data heldout.
@@ 4.7 DEL-Compose facilitates structured interpretation of the data
In Figure 8, we plot the attention distribution over the set of synthon representations of DEL-Compose. The top set of plots aggregate known pharmacophores, while the bottom set of plots delineate individual substructure sets. For CA-IX, we see that the attention probabilities are primarily focused on di-synthon $x_{bc}$, while for HRP, the attention probabilities are primarily distributed on the mono-synthon $x_{b}$. The sulfonamides are on the C position for CA-IX, while the electrophiles are on the B position for HRP, indicating that both models chose the highest weight to be placed on synthon embeddings incorporating the correct synthon position. Moreover, the weights on the important mono/di-synthons are higher for the molecules with the known pharmacophores versus other molecules in the library. Interestingly, the model for CA-IX chooses the di-synthon $x_{bc}$, instead of just the mono-synthon at the C position. A closer inspection of the enriched di-synthons reveals that the model also highly enriches synthons at the B position which contain sulfonyl chlorides. These synthons form sulfonamides in the tri-synthon structure, albeit internal ones, so the model also predicts that internal sulfonamides have some degree of affinity.
4.8 Applicability of DEL-Compose
DEL-Compose offers a structured way to characterize DEL data by decomposing the learned representations from a molecule’s synthon composition. Here, we have demonstrated the ability of DEL-Compose to capture important features for in-distribution molecules and to predict enrichments for molecules that align with known active substructures for the proteins in the datasets. Moreover, our model is generally applicable for learning useful representations for many other downstream tasks related to drug discovery. For instance, the representations learned from our model can be used to predict properties of molecules out-of-distribution, or to be used as a guide for generation of new molecules. This contrasts to other works that only try to derive signal from in-library molecules 9, 10, 27, 26. Moreover, we present a generative framework that captures the synthesize process of DEL molecules, which does not assume the availability of certain data such as reaction yields. We also demonstrate the viability of training such models without requiring fully enumerated molecule structures.
However, there are also limitations of our experiments here. Due to the nature of the data, our analysis only extends to tasks in which the important chemical substructures are localized within a single synthon, either at the internal B position, or the terminal C position. While we do show that our model attributes meaningful mono- and di-synthons as important, our model can require additional tools to learn powerful representations for more complex targets in order to better capture complex dependency structures. In the proposed DEL-Compose model, we do not consider the explicit chemical structure of di-synthons. However, we can add additional regularizers for the model to predict di-synthon and even tri-synthon structures so that the model can more effectively capture higher order structures. In our work, we also relied on substructure knowledge to split the data into synthon splits. But when such knowledge is unavailable, we can instead either use scaffold splits or cluster the molecules using molecular fingerprints and create splits within each cluster.
@@ 6 Data and Software Availability
Software: we use RDKit (version 2020.09.128 to parse our molecules and generate Morgan Fingerprints. The details to replicate our model is in the Methods section, and further training details are provided in the Experiments section.
{suppinfo}
6.1 CA-IX Enriched Synthons
From our analysis of our model’s attention module, we see that for CA-IX, the model assigns the highest weight to the di-synthon BC, instead of just mono-synthon C, where the terminal benzene-sulfonamides strucutres are located. When we look at which synthons at the B position are enriched, we see that this contains a list of sulfonyl chlorides (see Figure 9). Full molecules with these synthons have internal sulfonamides, so it is reasonable that the model would upweigh the molecules with these substructures. Interestingly, the internal sulfonamide is not actually apparently when only considering the synthons at the B position alone.
Figure 9: Top enriched synthons at the B position for CA-IX and example full molecules with these synthons (highlighted in blue).
We also see that there are multiple synthons enriched in only the matrix for CA-IX, we have included them in Figure 10 for reference. Our models predict enrichment in the matrix for molecules containing these synthons, but not the enrichment for the protein.
Figure 10: Top enriched synthons at the B position in the matrix data for CA-IX. These molecules were predicted to be enriched in the matrix but not the target by our models.